OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. x + 2y - z = 7 4x - y + 3z = -2 2x + 2y - z = 9 (2, 1, 3) (2, 1, -3) (-2, 1, -3) (-2, -1, 3)

OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. x + 2y - z = 7 4x - y + 3z = -2 2x + 2y - z = 9 Choices (2, 1, 3) (2, 1, -3) (-2, 1, -3) (-2, -1, 3)

OpenStudy (anonymous):

anyone here.

OpenStudy (whpalmer4):

Bueller? Bueller? :-)

OpenStudy (whpalmer4):

Do you know how to do the substitution method?

OpenStudy (anonymous):

Im just looking for the answer atm.

OpenStudy (whpalmer4):

Maybe someone else will help you, then.

OpenStudy (anonymous):

ok lol I am busy with other homework note my late replies.

OpenStudy (anonymous):

ANYONE HELP!